# Beyond Grid-Locked Voxels: Neural Response Functions for Continuous Brain Encoding

### v1

@@ Abstract
Neural encoding models aim to predict fMRI-measured brain responses to natural images. fMRI data is acquired as a 3D volume of voxels, where each voxel has a defined spatial location in the brain. However, conventional encoding models often flatten this volume into a 1D vector and treat voxel responses as independent outputs. This removes spatial context, discards anatomical information, and ties each model to a subject-specific voxel grid. We introduce the NRF Neural Response Function, a framework that models fMRI activity as a continuous function over anatomical space rather than a flat vector of voxels. NRF represents brain activity as a continuous implicit function: given an image and a spatial coordinate $(x,y,z)$ in standardized MNI space, the model predicts the response at that location. This formulation decouples predictions from the training grid, supports querying at arbitrary spatial resolutions, and enables resolution-agnostic analyses. By grounding the model in anatomical space, NRF exploits two key properties of brain responses: (1) local smoothness—neighboring voxels exhibit similar response patterns; modeling responses continuously captures these correlations and improves data efficiency, and (2) cross-subject alignment—MNI coordinates unify data across individuals, allowing a model pretrained on one subject to be fine-tuned on new subjects. In experiments, NRF outperformed baseline models in both intrasubject encoding and cross-subject adaptation. Achieving high performance while reducing the data size needed by orders of magnitude. To our knowledge, NRF is the first anatomically aware encoding model to move beyond flattened voxels, learning a continuous mapping from images to brain responses in 3D space.
Figure 1: Overview of NRF. Top: Individual-subject NRF. Brain responses are modeled as a continuous function of both image features and anatomical coordinates in MNI space. NRF learns how to map to voxel responses while capturing correlations between neighboring voxels through anatomical coordinates. Bottom: New subject adaptation. 1) For novel subjects, the learned representation could be transferred by fine-tuning pretrained NRFs with limited data. 2) Predictions from multiple finetuned base models are combined via voxelwise ensembling to capture individual variability. NRF thus moves beyond grid-locked voxel models, offering a continuous, anatomically grounded representation that enables both data-efficient single-subject encoding and flexible cross-subject transfer.
@@ 1 Introduction
A major goal in computational neuroscience is to understand how the human brain maps visual stimuli into neural activity. Neural encoding models aim to address this by predicting neural responses-typically measured by fMRI—from visual stimuli. These models offer powerful tools for analyzing high-dimensional brain data and probing the representations encoded in the visual Downing et al. (2001); Epstein & Kanwisher (1998); Gu et al. (2022); Heeger & Ress (2002); Kanwisher et al. (1997); Naselaris et al. (2011); Huth et al. (2012).
@@
However, the real-world utility of current neural encoding models remains limited. Current neural encoding models represent fMRI responses as a 1D vector in $R^{n}$, where n is a subject-specific voxel count. Naselaris et al. (2015)St-Yves & Naselaris (2018)Wang et al. (2023)Yamins et al. (2014) This discrete formulation has two critical limitations: 1) Ignoring 3D structure. By flattening fMRI volumes into 1D vectors, conventional models discard spatial information. This removes local context: anatomically adjacent voxels, which are often functionally correlated, are instead treated as independent outputs. 2)Subject specific. Each model is tied to the voxel grid of a single subject, making it non-transferable across individuals. Because voxel counts differ across brains, the output dimensionality of conventional models—and therefore their architecture—is tied to a single subject. As a result knowledge learned from one individual cannot be directly transferred to another, forcing each new subject to require training a separate model from scratch.
@@
As a result, models fail to exploit the 3D geometry of brain activity, wasting statistical power and requiring more data to learn accurate mappings. These issues are particularly problematic in realistic applications, where data are scarce. Unlike large-scale efforts such as NSD Allen et al. (2022), which collected tens of thousands of trials per subject over more than a year of scanning, most studies can only acquire a few hundred trials per subject due to cost and time constraints. Thus, the inefficiencies of discrete neural encoding models are amplified in real-world low-data regimes. In short, previous encoding models are grid-locked: they can only predict responses at the discrete sampling points they were trained on, for a single subject, and at one resolution. However, the human brain is a continuous 3D structure. Within each subject, neighboring voxels also exhibit similar response patterns, reflecting the spatial smoothness of neural activity. Despite individual variability, the visual cortex is highly conserved across people: areas such as the fusiform face area (FFA) and extrastriate body area (EBA) consistently respond to the same stimulus categories, and these regions align well across subjects when mapped into standardized anatomical templates such as MNI space. Ignoring this organization—both local smoothness and cross-subject correspondences — discards valuable structure that could enable more efficient learning and better generalization.
@@
  • Exploiting local smoothness. By conditioning predictions on anatomical coordinates, NRF incorporates the 3D spatial structure of fMRI data. This allows nearby voxels—often anatomically connected and functionally correlated—to share information, rather than being treated as independent outputs. As a result, NRF captures local smoothness in brain responses and achieves greater data efficiency.
  • Efficient new-subject adaptation. NRF grounds predictions in standardized MNI space, unifying responses across subjects in a shared coordinate system. This enables direct transfer: a model pretrained on one subject can be adapted to a new individual with only minimal fine-tuning. We further introduce a finetune–ensemble strategy that leverages multiple pretrained models to boost adaptation accuracy, reducing the need for extensive subject-specific data collection in real-world settings.
  • Resolution-agnostic predictions. By defining responses in continuous space, NRF decouples predictions from the voxel grid of the data model that is trained. It can be queried at arbitrary spatial coordinates, independent of the voxel size or sampling scheme used in data collection, and can seamlessly integrate data acquired at different resolutions. This flexibility opens the door to building more general-purpose brain models, moving closer to a functional digital twin of the brain.
Through experiments, we show that NRF - the first brain anatomy structure-grounded neural encoding model - offers a new direction for efficient and generalizable neural encoding.
@@ 2 Related Work
Neural Encoding models
fMRI encoding models have been extensively studied over the past two decades Mitchell et al. (2008); Huth et al. (2016); Gu et al. (2022); Tang et al. (2023); Kay et al. (2008); Güçlü & Van Gerven (2015); Naselaris et al. (2015). Most existing approaches treat fMRI data as discrete, formulating the problem as a regression task that maps image features to voxel-wise responses Naselaris et al. (2011); Han et al. (2019); Wang et al. (2023). In these models, fMRI responses are flattened as a 1-D vector, and each voxel is treated independently, ignoring the anatomical structure of the brain. As a result, they fail to capture the inherent local smoothness of fMRI responses, where neighboring voxels often exhibit correlated activity. To our knowledge, we are the first to develop an anatomically aware continuous encoding model, which leverages the 3D structure of the brain to improve data efficiency and generalization.
@@
Implicit neural representations have emerged as a powerful paradigm for modeling continuous signals in computer vision and graphics. Instead of storing data on fixed grids, INRs represent signals such as images Sitzmann et al. (2020) and 3D shapes Park et al. (2019); Mildenhall et al. (2021); Chen & Zhang (2019); Mescheder et al. (2019) as continuous functions parameterized by neural networks. A key advantage of this framework is its ability to capture fine-grained structure and support resolution-agnostic queries. Inspired by this line of work, we adopt a similar coordinate-based formulation for fMRI encoding. Unlike prior voxelwise models that discretize the brain into subject-specific grids, our approach treats brain responses as a continuous function over standardized anatomical coordinates. To our knowledge, NRF is the first attempt to bring the implicit representation framework to computational neuroscience, enabling anatomically aware, resolution-agnostic modeling of fMRI responses.
@@ 3 Method
3.1 Modeling Brain Response Mapping with implicit neural representation
@@
We instantiate $\Phi$ using a two-component design. The first component, $G$, the image feature extraction block. It extracts multi-scale features from the stimulus image $M$, capturing both low-level and high-level representations. These features are fused together to obtain a final image embedding $G(M)$. The second component is an implicit neural representation predictor $P$ that conditions on both $G(M)$ and the spatial coordinate $\mathbf{x}$. The coordinate is first encoded using Fourier features Tancik et al. (2020):
@@
The model is trained end-to-end with all components learned together, where the objective of the model is to correctly predict the voxel activation on each input image. Training batches are constructed from 32 randomly selected images, where for each image, we randomly sample 2000 voxels (out of 13000-15000 voxels) along with their corresponding fMRI activations for prediction. The model is trained using the Adam optimizer with a learning rate of 3e-3. For the loss function, we employ the same loss as in Beliy et al. (2019), a convex combination of mean square error and cosine similarity between the predicted response $\hat{r}$ and ground truth fMRI, $r$. The fMRI loss is defined as:
@@ 3.2 Cross Subject Transfer
NRF overcomes this limitation by being voxel-grid agnostic. Since responses are defined as a continuous function over standardized MNI space, subjects are naturally aligned in a shared anatomical coordinate system. This enables direct transfer: a model trained on one subject can be adapted to another without voxel-wise resampling, on the new subject’s coordinates and responses. Unlike classical voxelwise models—which rigidly tie the representation to a subject-specific grid—NRF learns a continuous, anatomically grounded representation that could flexibly generalize across individuals with only minimal data. To exploit this property, we adopt a two-step adaptation strategy:
@@
Voxelwise Ensemble. To further improve performance on new subjects, we perform a voxel-wise ensemble of the predictions from different finetuned models. Similar to the personalized ensemble approach in Gu et al. (2022), this strategy maximizes predictive performance while preserving inter-subject variability to improve model personalization. Specifically, for each voxel $v$, let $\{\hat{r}^{(1,i)}_{v},\hat{r}^{(2,i)}_{v},\dots,\hat{r}^{(K,i)}_{v}\}$ denote the predictions of $K$ finetuned base models for the $i$th image. We then learn voxel-specific weights $w_{v,k}$ (one per base model, for the $k$ th base model) and a bias $b_{v}$ by solving a least-squares regression on the limited new subject training data:
@@ 4 Experimental setup
We use the Natural Scene Dataset (NSD)Allen et al. (2022), which includes whole-brain 7T fMRI data from 8 subjects who viewed  10,000 natural scene images from the MS COCO dataset, repeated 1-3 times. The brain activations were computed using the GLMSingle algorithm Prince et al. (2022), and each voxel’s response value is normalized per session ($\mu=0,\sigma^{2}=1$). The brain activation to repeated images within a subject was averaged. The Neural response function(NRF) was trained using  9,000 unique images per subject, with around 1,000 images used for testing model accuracy via $R^{2}$. Since we are focusing on the visual cortex regions, we apply the official nsdgeneral region-of-interest (ROI) mask, which spans visual regions ranging from the early visual cortex to higher visual areas. Our evaluation focuses on Subj01, Subj02, Subj05, and Subj07 because these subjects completed all experiment sessions.
@@
To quantify prediction accuracy, we compute the voxel-wise Pearson correlation (Pearson) and voxel-wise mean square error (MSE) across all testing images.
Semantic-Level Metrics:
Following prior work Bao et al. (2025), we evaluate the semantic fidelity of predicted responses using MindEye2 Scotti et al. (2024), a pretrained fMRI-to-image decoder. The decoder reconstructs visual stimuli from predicted fMRI responses, which are then compared against the ground-truth stimuli presented during data collection. We employ a suite of image reconstruction metrics: PixCorr and SSIM quantify low-level visual fidelity, while Alex(2) and Alex(5) capture feature similarity at early and deeper layers of AlexNet. To assess semantic alignment, we further compute Incep, CLIP, Eff, and SwAV scores, which measure the correspondence between reconstructed and original images in higher-level representational spaces. Additional details on these metrics are provided in the Appendix.

### v2

@@ Abstract
Neural encoding models aim to predict fMRI-measured brain responses to natural images. fMRI data is acquired as a 3D volume of voxels, where each voxel has a defined spatial location in the brain. However, conventional encoding models often flatten this volume into a 1D vector and treat voxel responses as independent outputs. This removes spatial context, discards anatomical information, and ties each model to a subject-specific voxel grid. We introduce the NRF Neural Response Function, a framework that models fMRI activity as a continuous function over anatomical space rather than a flat vector of voxels. NRF represents brain activity as a continuous implicit function: given an image and a spatial coordinate $(x,y,z)$ in standardized MNI space, the model predicts the response at that location. This formulation decouples predictions from the training grid, supports querying at arbitrary spatial resolutions, and enables resolution-agnostic analyses. By grounding the model in anatomical space, NRF exploits two key properties of brain responses: (1) local smoothness—neighboring voxels exhibit similar response patterns; modeling responses continuously captures these correlations and improves data efficiency, and (2) cross-subject alignment—MNI coordinates unify data across individuals, allowing a model pretrained on one subject to be fine-tuned on new subjects. In experiments, NRF outperformed baseline models in both intrasubject encoding and cross-subject adaptation. Achieving high performance while reducing the data size needed by orders of magnitude. To our knowledge, NRF is the first anatomically aware encoding model to move beyond flattened voxels, learning a continuous mapping from images to brain responses in 3D space. Code and project site: https://github.com/haomiao8/NRF
Figure 1: Overview of NRF. Top: Individual-subject NRF. Brain responses are modeled as a continuous function of both image features and anatomical coordinates in MNI space. NRF maps image features to voxel responses while capturing correlations between neighboring voxels through shared anatomical coordinates. Bottom: New-subject adaptation. For a novel subject, we start from a pretrained individual-subject NRF (trained on the other subjects) and then fine-tune this model using only a small amount of data from the new subject. The figure shows both the pretrained and the fine-tuned NRFs to make explicit how the base model is adapted during transfer. Predictions from multiple fine-tuned base models are then combined via voxel-wise ensembling to better capture subject-specific variability. NRF thus moves beyond grid-locked voxel models, providing a continuous, anatomically grounded representation that supports both data-efficient single-subject encoding and flexible cross-subject transfer.
@@ 1 Introduction
A major goal in computational neuroscience is to understand how the human brain maps visual stimuli into neural activity. Neural encoding models aim to address this by predicting neural responses-typically measured by fMRI—from visual stimuli. These models offer powerful tools for analyzing high-dimensional brain data and probing the representations encoded in the visual system  (Downing et al., 2001; Epstein & Kanwisher, 1998; Gu et al., 2022; Heeger & Ress, 2002; Kanwisher et al., 1997; Naselaris et al., 2011; Huth et al., 2012) .
@@
However, the real-world utility of current neural encoding models remains limited. Current neural encoding models represent fMRI responses as a 1D vector in $\mathbb{R}^{n}$, where n is a subject-specific voxel count (Naselaris et al., 2015; St-Yves & Naselaris, 2018; Wang et al., 2023; Yamins et al., 2014). This discrete formulation has two critical limitations: 1) Ignoring 3D structure. By flattening fMRI volumes into 1D vectors, conventional models discard spatial information. This removes local context: anatomically adjacent voxels, which are often functionally correlated, are instead treated as independent outputs. 2) Subject specific. Each model is tied to the voxel grid of a single subject, making it non-transferable across individuals. Because voxel counts differ across brains, the output dimensionality of conventional models and their corresponding architectures are tied to a single subject. As a result, knowledge learned from one individual cannot be directly transferred to another, forcing each new subject to require training a separate model from scratch. Consequently, these models fail to exploit the 3D geometry of brain activity, wasting statistical power and requiring more data to learn accurate mappings.
@@
These issues are particularly problematic in realistic applications, where data are scarce. Unlike large-scale efforts such as NSD  (Allen et al., 2022), which collected tens of thousands of trials per subject over a year-long scanning period, most studies can only acquire a few hundred trials per subject due to cost and time constraints. Thus, the inefficiencies of discrete neural encoding models are amplified in real-world low-data regimes. In short, previous encoding models are grid-locked: they can only predict responses at the discrete sampling points they were trained on, for a single subject, and at one resolution.
However, the human brain is a continuous 3D structure. Within each subject, neighboring voxels also exhibit similar response patterns, reflecting the spatial smoothness of neural activity. Despite individual variability, the visual cortex is highly conserved across people :areas such as the fusiform face area (FFA) and extrastriate body area (EBA) consistently respond to the same stimulus categories, and these regions align well across subjects when mapped into standardized anatomical templates such as MNI space  (Heeger & Ress, 2002; Kanwisher et al., 1997; Downing et al., 2001; Epstein & Kanwisher, 1998; Heeger & Ress, 2002; Naselaris et al., 2011; Huth et al., 2012). Ignoring this organization—both local smoothness and cross-subject correspondences — discards valuable structure that could enable more efficient learning and better generalization.
@@
  • Exploiting local smoothness. By conditioning predictions on anatomical coordinates, NRF incorporates the 3D spatial structure of fMRI data. This allows nearby voxels that are anatomically connected and functionally correlated to share information instead of being treated as independent outputs. As a result, NRF captures local smoothness in brain responses and achieves greater data efficiency.
  • Efficient cross-subject adaptation. NRF grounds predictions in standardized MNI space, unifying responses across subjects in a shared coordinate system. This enables direct transfer: a model pretrained on one subject can be adapted to a new individual with only minimal fine-tuning. We further introduce a finetune–ensemble strategy that leverages multiple pretrained models to boost adaptation accuracy, reducing the need for extensive subject-specific data collection in real-world settings.
  • Resolution-agnostic modeling. By defining responses in continuous space, NRF decouples predictions from the specific voxel grid used during training. The model can be queried at arbitrary spatial coordinates, independent of the voxel size or sampling scheme used during data collection. This allows for the seamless integration of data acquired at varying resolutions and opens the door to building general-purpose brain models that move closer to a functional digital twin of the human visual system.
Through experiments,we demonstrate that NRF provides a novel anatomy-grounded framework for neural encoding, offering a new paradigm for efficient and generalizable brain modeling.
@@ 2 Related Work
Neural encoding models.
fMRI encoding models have been extensively studied over the past two decades  (Mitchell et al., 2008; Huth et al., 2016; Gu et al., 2022; Tang et al., 2023; Kay et al., 2008; Güçlü & Van Gerven, 2015; Naselaris et al., 2015). Most existing approaches treat fMRI data as discrete, formulating the problem as a regression task that maps image features to voxel-wise responses  (Naselaris et al., 2011; Han et al., 2019; Wang et al., 2023).Recent work has also explored stronger transformer-based encoders  (Adeli et al., 2025; Beliy et al., 2024). In these models, fMRI responses are flattened as a 1-D vector, and each voxel is treated independently, ignoring the 3D anatomical structure of the brain. While some approaches incorporate spatial priors, such as fitting multi-parameter models for spatial frequency mapping  (Broderick et al., 2022) or using Bayesian templates to warp retinotopic maps  (Benson & Winawer, 2018). However, general encoding frameworks largely fail to exploit the local smoothness and cross-subject correspondences inherent in brain activity. To our knowledge, NRF is one of the first anatomically aware encoding models that formulates image to fMRI prediction mapping as a continuous function over 3D brain space, leveraging anatomical structure to improve data efficiency and generalization.
Neural decoding models.
A parallel line of work focuses on reconstructing or decoding visual stimuli from fMRI signals. Early approaches  (Scotti et al., 2024) typically utilize subject-specific MLPs that cannot naturally generalize across individuals due to varying voxel counts and layouts. Recent methods address this via adaptive pooling for cross-subject decoding  (Wang et al., 2024) or by incorporating voxel (x,y,z) coordinates into attention mechanisms via positional encodings  (Qiu et al., 2025). While these methods highlight the importance of anatomical alignment, they pursue a fundamentally different goal: decoding information from measured, discrete voxel grids. Even with spatial cues, their operations remain tied to the specific voxels acquired. In contrast, we address the encoding problem by treating the brain as a continuous 3D structure. Rather than using coordinates as auxiliary features for discrete voxels, NRF models neural activity as a continuous function in brain space. This formulation enables resolution-agnostic querying at arbitrary positions and facilitates seamless cross-subject adaptation within a unified anatomical coordinate system.
Brain semantic mapping.
A rich line of research has mapped the semantic organization of the cortex using voxel-wise encoding models  (Huth et al., 2016; Deniz et al., 2019). These studies typically train independent encoders on flattened 1D response vectors and project learned selectivities onto the cortical surface post hoc for visualization. While this pipeline uses 3D spatial information during analysis, the encoding models themselves do not incorporate anatomical structure during training; each voxel is modeled independently, ignoring its physical location. NRF differs by injecting anatomical structure directly into the learning process. Rather than using coordinates solely for downstream visualization, NRF conditions the encoding function on 3D $(x,y,z)$ coordinates and learns a continuous neural field over standardized brain space. This allows the model to operationalize the spatial principles revealed in prior studies, such as the fact that nearby voxels exhibit smoothly varying semantic selectivity. By modeling neural activity as a spatially smooth function rather than a set of independent points, NRF achieves greater data efficiency and enables flexible cross-subject adaptation without the need to resample volumes onto a shared grid.
@@
Implicit neural representations(INR) have emerged as a powerful paradigm for modeling continuous signals in computer vision and graphics. Instead of storing data on fixed grids, INRs represent signals such as images  (Sitzmann et al., 2020) and 3D shapes  (Park et al., 2019; Mildenhall et al., 2021; Chen & Zhang, 2019; Mescheder et al., 2019) as continuous functions parameterized by neural networks. A key advantage of this framework is its ability to capture fine-grained structure and support resolution-agnostic queries. Inspired by this line of work, we adopt a similar coordinate-based formulation for fMRI encoding. Unlike prior voxel-wise models that discretize the brain into subject-specific grids, our approach treats brain responses as a continuous function over standardized anatomical coordinates. To our knowledge, NRF is the first attempt to bring the implicit representation framework to computational neuroscience, enabling anatomically aware, resolution-agnostic modeling of fMRI responses.
@@ 3 Method
3.1 Modeling Neural Response as implicit neural representation
@@
We instantiate $\Phi$ using a two-component design. The first component, $G$, the image feature extraction block. It extracts multi-scale features from the stimulus image $M$, capturing both low-level and high-level representations. These features are fused together to obtain a final image embedding $G(M)$. The second component is an implicit neural representation predictor $P$ that conditions on both $G(M)$ and the spatial coordinate $\mathbf{x}$. The coordinate is first encoded using Fourier features  (Tancik et al., 2020):
@@
The model is trained end-to-end with all components learned together, where the objective of the model is to correctly predict the voxel activation on each input image. Training batches are constructed from 32 randomly selected images, where for each image, we randomly sample 2000 voxels (out of 13000-15000 voxels) along with their corresponding fMRI activations for prediction. The model is trained using the Adam optimizer with a learning rate of 3e-3. For the loss function, we employ the same loss as in  (Beliy et al., 2019), a convex combination of mean square error and cosine similarity between the predicted response $\hat{r}$ and ground truth fMRI, $r$. The fMRI loss is defined as:
@@ 3.2 Cross Subject Transfer
NRF overcomes this limitation by being voxel-grid agnostic. Since responses are defined as a continuous function over standardized MNI space, subjects are naturally aligned in a shared anatomical coordinate system. This enables direct transfer: a model trained on one subject can be adapted to another without voxel-wise resampling, on the new subject’s coordinates and responses. Unlike classical voxel-wise models—which rigidly tie the representation to a subject-specific grid—NRF learns a continuous, anatomically grounded representation that could flexibly generalize across individuals with only minimal data. To exploit this property, we adopt a two-step adaptation strategy:
@@
voxel-wise Ensemble. To further improve performance on new subjects, we perform a voxel-wise ensemble of the predictions from different finetuned models. Similar to the personalized ensemble approach in  (Gu et al., 2022), this strategy maximizes predictive performance while preserving inter-subject variability to improve model personalization. Specifically, for each voxel $v$, let $\{\hat{r}^{(1,i)}_{v},\hat{r}^{(2,i)}_{v},\dots,\hat{r}^{(K,i)}_{v}\}$ denote the predictions of $K$ finetuned base models for the $i$th image. We then learn voxel-specific weights $w_{v,k}$ (one per base model, for the $k$ th base model) and a bias $b_{v}$ by solving a least-squares regression on the limited new subject training data:
@@ 4 Experimental setup
We use the Natural Scene Dataset (NSD) (Allen et al., 2022), which includes whole-brain 7T fMRI data from 8 subjects who viewed  10,000 natural scene images from the MS COCO dataset, repeated 1-3 times. The brain activations were computed using the GLMSingle algorithm  (Prince et al., 2022), and each voxel’s response value is normalized per session ($\mu=0,\sigma^{2}=1$). The brain activation to repeated images within a subject was averaged. The Neural response function(NRF) was trained using  9,000 unique images per subject, with around 1,000 images used for testing model accuracy via voxel-wise pearson correlation ($r$). Since we are focusing on the visual cortex regions, we apply the official nsdgeneral region-of-interest (ROI) mask, which spans visual regions ranging from the early visual cortex to higher visual areas. Our evaluation focuses on Subj01, Subj02, Subj05, and Subj07 because these subjects completed all experiment sessions.
@@
To quantify prediction accuracy, we compute the voxel-wise Pearson correlation ($r$) and voxel-wise mean square error (MSE) across all testing images.
Semantic-Level Metrics.
Following prior work  (Bao et al., 2025), we evaluate the semantic fidelity of predicted responses using MindEye2  (Scotti et al., 2024), a pretrained fMRI-to-image decoder. Given an input image, we first predict its fMRI response using NRF; the predicted response is then fed into MindEye2, which reconstructs the corresponding visual stimulus. We compare these reconstructions against the ground-truth images presented during data collection.We employ a suite of image reconstruction metrics. PixCorr and SSIM quantify low-level visual fidelity, while Alex(2) and Alex(5) measure similarity in early and deeper layers of AlexNet. To evaluate higher-level semantic alignment, we compute Incep, CLIP, Eff, and SwAV scores, which assess the representational correspondence between reconstructed and original images in diverse semantic embedding spaces. Additional details are provided in Appendix  A.3.